Solve the problem. Suppose that replacement times for washing machines are normally distributed with a mean of 9.3 years and a standard deviation of 1.1 years. Find the probability that 70 randomly selected washing machines will have a mean replacement time less than 9.1 years. Write your answer as a decimal rounded to 4 places.
Answer #1

\mu = 9.3

\sigma = 1.1

n = 70

SE = \sigma /\sqrt{n}

= 1.1/\sqrt{70} = 0.1315

To find P(\bar{x}< 9.1):

Z = (\bar{x} - \mu )/SE

=(9.1 - 9.3)/0.1315 = - 1.5212

Table of Area Under Standard Normal Curve gives area = 0.4357

So

P(\bar{x}<9.1) = 0.5 - 0.4357 = 0.0643

So,

Answer is:

0.0643